# Expert Study on Interpretable Machine Learning Models with Missing Data

### v1

@@ Abstract
Inherently interpretable machine learning (IML) models provide valuable insights for clinical decision-making but face challenges when features have missing values. Classical solutions like imputation or excluding incomplete records are often unsuitable in applications where values are missing at test time. In this work, we conducted a survey with 71 clinicians from 29 trauma centers across France, including 20 complete responses to study the interaction between medical professionals and IML applied to data with missing values. This provided valuable insights into how missing data is interpreted in clinical machine learning. We used the prediction of hemorrhagic shock as a concrete example to gauge the willingness and readiness of the participants to adopt IML models from three classes of methods. Our findings show that, while clinicians value interpretability and are familiar with common IML methods, classical imputation techniques often misalign with their intuition, and that models that natively handle missing values are preferred. These results emphasize the need to integrate clinical intuition into future IML models for better human-computer interaction.
keywords:
Interpretable machine learning, missing values, evaluation with clinicians
@@ 1 Introduction
Interpretable machine learning (IML) models have emerged as vital tools in high-stakes decision-making areas, such as healthcare (Ustun and Rudin, 2019a), sustainability (Jin et al., 2022), and criminal justice (Wang et al., 2023; Zhang et al., 2022), where transparency and accountability are paramount. Inherently interpretable models, such as linear or rule-based models, are well studied due to their apparent simplicity (Molnar, 2020; Rudin, 2019). While these models prioritize transparency, more advanced methods like interpretable decision sets (Rudin et al., 2022), checklists (Zhang et al., 2022) or falling rule lists (Wang and Rudin, 2015), effectively balance interpretability with predictive power, making them ideal for complex, real-world applications (Rudin et al., 2022). These models not only provide accurate predictions but also foster a deeper understanding of the model’s inner workings, enabling domain experts to make informed decisions (Molnar, 2020). Despite advancements in IML, aligning these models with clinical needs remains challenging (Rolnick et al., 2024). A key issue in healthcare applications is the prevalence of missing values both during training and in deployment, which can arise from incomplete patient records or errors during data collection (Salgado et al., 2016). Missing values are often addressed using imputation methods (Little and Rubin, 2019), which is general and powerful, but not always optimal under test-time missingness (Le Morvan et al., 2020b). Other methods like complete case analysis are inefficient (Penny and Atkinson, 2012), biased under non-uniform missingness, and are inapplicable to test-time cases with missing values. On the other hand, approaches that use simple imputations and missingness indicators (Le Morvan et al., 2020a) complicate interpretability. In short, missing values tend to make interpreting model predictions more difficult, potentially leading to issues during deployment.
To bridge the gap between IML model development and clinical use with test-time missingness, understanding clinicians’ needs is essential. A human-centered approach (Doshi-Velez and Kim, 2017) can guide model design to align with clinical expectations and improve healthcare reliability. We surveyed 71 clinicians to: 1) understand their interactions with predictive models, especially with missing values; 2) evaluate IML preferences aligned with clinical intuition; and 3) recommend improvements for handling missing data.
2 Methods
Participant recruitment
We conducted an online survey via LimeSurvey with 71 clinicians in July and August 2024 to evaluate their understanding of IML models and their effectiveness in handling missing data. Participants were recruited via email from TraumaBase network hospitals, and informed consent was obtained from all.
Experimental design
The survey, developed in English with input from trauma physicians and ML researchers, was translated into French by native speakers to ensure accuracy and cultural relevance. It included 41 questions on demographics, structured choice, and ranking tasks, Likert-scale statements, and open-ended questions, with both language versions provided in Appendix A. We conducted a pilot with five clinicians and incorporated their feedback. The survey aimed to understand clinicians’ interaction with IML when facing missing values at test time, using hemorrhagic shock prediction as an ICU decision-making example (Dutton et al., 2010) (details in Appendix A.2). The IML models included well-known interpretable classes such as risk scores (Ustun and Rudin, 2019b), linear model coefficients (Hastie et al., 2009), and decision trees (Breiman, 2017), all with comparable predictive performance (Molnar, 2020; Stiglic et al., 2020), as predictive performance was not the focus here. We also examined how different imputation methods influenced clinicians’ decisions. Presented imputation methods included native methods (e.g., XGBoost (Chen and Guestrin, 2016)) without imputation, implicit imputation allowing clinicians to adjust based on other feature values and judgment, explicit methods like mean or zero imputation using statistical replacements, and complex approaches such as MICE (Van Buuren and Groothuis-Oudshoorn, 2011) representing black-box methods. Decision trees handle missing data through surrogate splits. Figure LABEL:fig:IML_survey illustrates the survey interface, showing a case with one missing variable.
@@ 2 Methods
fig:IML_survey Refer to caption
@@
Data analysis
We analyzed multiple-choice responses using descriptive statistics, while Likert scale and categorical data were examined with factor analysis for mixed data (FAMD) using the FactoMineR package in R (Lê et al., 2008) to identify patterns and group perspectives. Free-text responses were analyzed for central tendency and variability, with selected French responses translated into English. Figure LABEL:fig:survey_example in the Appendix shows the experimental setup.
3 Results
3.1 Sample characteristics
The survey was administered to 214 clinicians across 29 trauma centers in France, part of the TraumaBase network, which collects data on over 50,000 trauma cases nationwide. We received 55 responses, with 20 (36.4%) fully completed and 35 (63.6%) partially completed. All responses were analyzed, so answer counts vary by question. The average age was 41.2 years, with a gender distribution of 22 females and 33 males; most participants were clinical doctors (83%) in Anesthesiology/Intensive Care (84%) with medical degrees (84%). Participant demographics are in Table LABEL:app:tab:stats_participants. Our FAMD analysis identified four clusters of clinicians based on AI/ML attitudes, familiarity with IML, and missing value handling preferences (analysis in Appendix A.4). The most prominent cluster (25%) comprises frequent AI users, highly familiar with IML models, and prefer implicit imputation over black-box methods.
Current practices with missing values
Clinicians use various strategies to manage missing data in clinical scores, primarily relying on clinical intuition or other available clinical information. Many mentioned using plausible values based on clinical context or the patient’s history to estimate missing values. Some prefer worst-case imputation to avoid underestimating risk, though one noted it may lead to overtraining and resource inefficiency. Others choose to partially calculate or omit scores if key data is missing, feeling that imputation could ”skew the score concept.” One clinician also suggested using alternatives, like thromboelastography when prothrombin time is unavailable.
3.2 Evaluation of IML with missing values
Our survey focuses on evaluating how clinicians interact with tools, investigating how different predictive models handle (or natively handle) missing values, and how clinicians interpret these models to make informed decisions. We also examine what clinicians find lacking in current IML approaches and what improvements are needed to make these models more useful for clinical decision-making.
Understanding approaches to handling missingness in prediction tasks
To understand clinicians’ perspectives on common imputation methods, we first provided average hemoglobin values and tested reactions to zero-imputed data (Table LABEL:tab:Imputation_ranking). Most clinicians reported that average values were irrelevant for decision-making, emphasizing the need for individualized consideration, as one explained: “Each patient is unique, and I don’t see the relevance of giving an average value.” Additionally, some clinicians expressed concern that mean imputation might underestimate patient risk, potentially minimizing preparedness for emergency scenarios. When presented with a zero-imputed value, many clinicians indicated they would ignore the feature, though some noted the importance of clearly marking any imputation along with associated uncertainty. For decision trees, using surrogate values was explored, but this method was less familiar to clinicians and caused some confusion. However, many clinicians noted that its usefulness depends on clearly indicating where the substitution occurs. More complex methods as MICE which frequently used and cited in medical journals (e.g.,  Austin et al. (2021); Hegde et al. (2019); Luo (2022)) as an imputation tool for clinical data. While clinicians recognize MICE for its effectiveness in imputing missing values, it does not necessarily enhance predictive performance. MICE is designed for inference, not primarily for supervised learning. In our survey, black-box methods like MICE were selected in less than 10% of cases (6/51), indicating that clinicians hesitate to use methods they cannot interpret easily. MICE’s low preference may also stem from it not being commonly taught to physicians.
Clinicians are confident but not always correct under missing values
Clinicians showed high confidence (9-10 out of 20 answers) in explaining risk scores and decision trees, even with missing values, and interpreted the model outcome correctly. For logistic regression, however, confidence was lower and accompanied by uncertainty; only 6 participants were correct, with 4 out of 6 correct among those expressing confidence in its usefulness with missing values.
Requirements to support interaction between (future) IML & clinicans under missing values
Four out of 12 clinicians emphasized the need for clear guidelines and transparency in handling missing data, with one suggesting the presentation of multiple imputed values to allow intuitive selection along with communicating uncertainty. The reliability of models for managing missing data were seen as essential, with two clinicians stressing that ML conclusions should be accessible to non-technical users, recommending external validation and visual aids to clarify the impact of imputation. Clinicians underscored the importance of integrating clinical intuition with ML outputs, emphasizing that “the human element and intuition based on experience and the CLINICAL EXAMINATION OF THE PATIENT!” should complement, not replace, the model. Their preference for implicit over explicit imputation methods reflects a focus on transparency and alignment with clinical reasoning, as implicit methods enhance situational awareness.
4 Discussion
Context
While Janssen et al. (2010) argues that simple methods for handling missing data can lead to misleading results, others advocate for multiple imputation techniques as they enhance validity in clinical research (Austin et al., 2021). Following Tonekaboni et al. (2019), who emphasized evaluating clinical explainability in high-stakes settings, IML models must undergo rigorous validation to assess performance, interpretability, and legal implications (Wiens et al., 2019). Clinicians can build trust in IML models by understanding them and through positive experiences where model outputs align with domain knowledge, enhancing appreciation for the model’s utility while ensuring accountability for patient outcomes (Kelly et al., 2019). Randomized controlled trials (RCTs) can validate models, but are costly, time-consuming, and may lack generalizability across diverse clinical settings (Shortliffe and Sepúlveda, 2018). Successful deployment involves robust model development, presenting uncertainties (Doshi-Velez and Kim, 2017), and training clinicians on the impact of missing values to integrate these tools into workflows (Topol, 2019; Maddox et al., 2019).
Limitations
Our survey focused on ICU and trauma specialists with foundational knowledge of clinical ML systems. However, it may reflect a self-selection bias due to the nature of participant recruitment. Focused on IML models for prediction with missing values, the findings may not directly apply to other paradigms, like reinforcement learning, or to settings like outpatient care. Only models from three IML classes were explored and the scope was limited to clinicians in France, highlighting the need for broader research across different countries and healthcare systems to generalize and externally validate the results.
Development of IML models with missing values: Interpretation of clinical needs
Our results reveal that current models often fall short of supporting clinical decision-making when faced with missing data at test time. To improve adoption by healthcare professionals, future IML models should better align with clinical intuition and incorporate clinical experience when handling missing data.
While imputation methods like zero, mean, or black-box imputation may facilitate model training, they are less favored for test-time use. Worst-case imputation, although cautiously applied by some clinicians, can introduce statistical bias, skew predictions, and misrepresent patient risk. Clinicians prefer that imputation, if necessary, incorporate patient-specific data and clinical experience, rather than generic values like zeros or cohort means. Moreover, any imputation used should be transparent, as clinicians bear responsibility for decisions impacting patient outcomes. Future methods should either avoid reliance on imputation or employ more transparent imputation strategies that are suitable both during training and in use. Models that natively support missing values, like those by Chen and Guestrin (2016) and Stempfle and Johansson (2024), eliminate or reduce the need for imputation during training or testing. Alternatively, more transparent imputation methods can enhance trust. For example, recent work by Chen et al. (2023) on explainable boosting machines provides insights into missing data mechanisms, helping to understand causes and mitigate risks. As mentioned by several respondents, future models should also respect clinical reasoning by integrating patient history and the sequential nature of medical information. Current IML methods often overlook this aspect, making alternatives like reinforcement learning potentially more suited to capturing the complexity of clinical decision-making.
Conclusion
The results reveal that clinicians value IML models but find traditional approaches like imputation inadequate when data is missing at test time. They prefer models that can natively handle missing values, as imputation often conflicts with clinical intuition. This underscores the need for ML models aligned with clinical needs, particularly in high-stakes areas like hemorrhagic shock prediction. The study highlights clinicians’ preferences for advanced methods to enhance decision-making and patient outcomes with more reliable, interpretable tools.

### v2

@@ Abstract
Inherently interpretable machine learning (IML) models offer valuable support for clinical decision-making but face challenges when features contain missing values. Traditional approaches, such as imputation or discarding incomplete records, are often impractical in scenarios where data is missing at test time. We surveyed 55 clinicians from 29 French trauma centers, collecting 20 complete responses to study their interaction with three IML models in a real-world clinical setting for predicting hemorrhagic shock with missing values. Our findings reveal that while clinicians recognize the value of interpretability and are familiar with common IML approaches, traditional imputation techniques often conflict with their intuition. Instead of imputing unobserved values, they rely on observed features combined with medical intuition and experience. As a result, methods that natively handle missing values are preferred. These findings underscore the need to integrate clinical reasoning into future IML models to enhance human-computer interaction.
@@ 1 Introduction
Missing values in healthcare data pose substantial challenges in predictive modeling, requiring effective methods to address them (Austin et al., 2021; Wells et al., 2013). These missing values occur in incomplete medical records and often arise intentionally, due to specific testing protocols, or unintentionally, due to time constraints or data collection errors (Salgado et al., 2016). The resulting gaps can significantly impact the reliability and applicability of predictive models, particularly at ”test time” when they are used for decision-making. Interpretable machine learning (IML) models have become essential in high-stakes decision-making areas such as healthcare (Ustun and Rudin, 2019a), sustainability (Jin et al., 2022), and criminal justice (Wang et al., 2023; Zhang et al., 2022), where transparency and accountability are critical. These models offer insights into decision processes, ensuring that predictions can be understood and trusted by domain experts. IMLs, such as decision trees, linear models, and rule-based models, are extensively studied for their intuitive structure and ease of interpretation (Molnar, 2020; Rudin, 2019). They are widely used in tasks like classification, regression, and risk scoring (Fürnkranz et al., 2012; Margot and Luta, 2021). As with most ML models, IML models to date rely on complete inputs, making them particularly vulnerable to missing data. In retrospective studies, a typical solution is complete-case analysis, where cases with missing values are excluded (Little and Rubin, 2019). While effective for analysis, this approach is unsuitable at test time, as it can leave many patients without predictions. An alternative is the impute-then-regress strategy which imputes missing values before prediction, ensuring consistency between training and deployment tasks (Rubin, 1976). However, impute-then-regress is often biased when the missingness of predictive features depends on unobserved variables. Moreover, complex imputation methods can reduce interpretability, complicating deployment and decision-making. Despite extensive research on handling missing values (Van Buuren and Groothuis-Oudshoorn, 2011; Le Morvan et al., 2020; Chen et al., 2023; Austin et al., 2021; Luo, 2022; Stempfle and Johansson, 2024), little is known about how medical domain experts interact with missing values when using prediction models.
@@
fig:objectives Refer to caption
Figure 1: Overview of study objectives, key survey question types, and findings on best practices for clinical interpretable machine learning with missing values at test time.
In this work, we survey clinicians to gain qualitative insights into their handling of missing values using IMLs, providing a real-world perspective on decision-making. Our evaluation is structured around three key objectives (shown in Figure LABEL:fig:objectives). First, we aim to understand clinicians’ attitudes toward artificial intelligence/machine learning (AI/ML) in healthcare and their current practices for clinical decision-making with missing values (Objective I). Next, we present a use case aimed at identifying common interpretations and usages of three different IML approaches and preferences for strategies to handle missing values (Objective II). Finally, we define requirements for future IML models that handle missing values at test time, ensuring they align with clinician preferences and can be seamlessly integrated into clinical workflows (Objective III).
Our survey focuses on a real-world scenario in handling missing values for interpretable clinical prediction. This complements and contrasts with the rich literature on quantitative and theoretical comparisons between, e.g., different imputation methods. We emphasize domain expertise through direct quotes and publish our full questionnaire, enabling collaborative research on IML and missing values to identify best practices across healthcare systems.
Our main contributions are as follows: 1) We clustered the survey responses of 55 survey participants to analyze clinicians’ attitudes toward AI/ML and their practices for handling missing values, providing an overview of existing workflows. 2) We assessed clinicians’ use of three different IML approaches in scenarios with missing values, analyzing their reasoning, preferences, and challenges in leveraging these models for decision-making. 3) Based on clinician feedback, we formulated guidelines for designing IML models that handle missing values natively, ensuring better alignment with clinical workflows and facilitating real-world adoption.
Our findings highlight a gap between missing value handling in machine learning research, which often employs overly simplistic or too complex imputation methods, and clinical decision-making, where clinicians prioritize observed data and medical intuition at the time of decision-making. Future approaches should focus on reducing reliance on imputation by adopting more interpretable strategies, with models that natively handle missing values.
2 Survey methodology
Participant recruitment.
We conducted an online survey using LimeSurvey  [^1] https://www.limesurvey.org in July–August 2024 to assess clinicians’ understanding of IML models and handling of missing values in decision-making. The survey involved 214 clinicians from 29 trauma centers across France, all part of the TraumaBase  [^2] https://www.traumabase.eu/en_US network, which collects data on over 50,000 trauma cases annually. Responses were assumed to be largely independent due to the geographical spread of centers. To ensure clinical relevance, we focused on predicting hemorrhagic shock (Dutton et al., 2010), a critical ICU scenario where missing information is common. In trauma ICUs, data is often incomplete due to lab results being unavailable before critical decisions are needed. Trauma ICU clinicians, accustomed to clinical scores that align with IML frameworks, are well-suited to evaluate IML models under test-time missingness. To boost response rates, we focused on a single use case and kept the survey within 20–25 minutes (details in Appendix A). Participants provided informed consent before participation.
Survey development.
The survey was developed with the input of trauma physicians and machine learning researchers to ensure its relevance to both clinical and ML communities. To maintain linguistic accuracy and cultural relevance, it was provided in English and French by native speakers. The full survey in both languages is included in Appendix B.
Survey design.
The survey consisted of 41 questions covering demographics, multiple-choice questions, ranking tasks, Likert-scale statements, and open-ended questions, aligned with the survey’s three objectives (Figure LABEL:fig:objectives). The first part of the survey focuses on understanding clinicians’ perspectives on AI/ML models within healthcare and their current approaches to handling missing values during decision-making (Objective I). It includes multiple-choice and Likert-scale questions to capture detailed insights and measure attitudes consistently. Questions on current practices referenced trauma-related risk scores, such as the Red Flag (Hamada et al., 2018) or ABC score (Nunez et al., 2009), ensuring familiarity and serving as a warm-up by linking familiar tools to the challenges of missing values. The second part of the survey examines a specific use case (Objective II), where clinicians predicted hemorrhagic shock for a patient with missing values using various IML models. It assesses how clinicians engaged with model outputs for decision-making, focusing on interpretability and handling incomplete data. Multiple-choice and ranking tasks uncovered preferences and priorities when interpreting IML models (details in the next paragraph). Finally, the third part comprises open-ended questions, allowing clinicians to share insights, contributing to guidelines for future IML research (Objective III).
IML and missing values in a use case.
The second part of the survey presented a patient case with a missing value and predictions by three IML models (Figure LABEL:fig:IML_survey). Participants were asked how they would handle the missing value in their decision-making process. This allows us to investigate whether clinicians would rely on additional information, such as a subpopulation’s average feature value, assign zero for the missing value, or apply their domain knowledge to estimate it. Additionally, we asked for their preferences in imputation strategies for different IML methods to better understand their approach to decision-making.
The survey asked questions about three well-established IML models: risk scores (Ustun and Rudin, 2019b), linear models (Hastie et al., 2009), and decision trees (Breiman, 2017). Since our primary interest was interpretability rather than predictive performance, we ensured all models had comparable accuracy (Molnar, 2020; Stiglic et al., 2020).
Handling missing values is challenging for these models, as they lack native support and often rely on imputation for predictions. Linear models require explicit handling of missing values since each feature contributes to the prediction based on its learned coefficient. If a feature is missing, its contribution is unknown unless imputed, meaning the model implicitly assumes the imputation strategy is correct. Rule-based risk scores require all features referenced in a rule to be available; otherwise, rule activation may be uncertain or impossible. Decision trees struggle with missing values due to their hierarchical structure, where each node depends on feature availability. If a feature is missing, the tree may follow a default branch (e.g., XGBoost Chen and Guestrin (2016)) or use imputed values, both potentially introducing bias. While XGBoost learns the default direction during training, this approach lacks interpretability at test time. Alternatively, decision trees can use surrogate splits (Feelders, 1999), identifying alternative features that mimic the primary split, enabling predictions without explicit imputation.
We use methods feasible for handling test-time missingness: imputation and native approaches:
  • Native methods: These are model-inherent mechanisms for handling missing values, such as surrogate splits in decision trees, which rely on alternative features to make predictions (Feelders, 1999).
  • Implicit imputation: Clinicians were presented with missing values without automated imputation and asked to estimate them using available features and their clinical judgment.
  • Explicit imputation: These included simple statistical replacements, such as mean or zero imputation (Rubin, 1976).
  • Black-box methods: More complex imputation techniques, such as Multiple Imputation by Chained Equations (MICE) (Van Buuren and Groothuis-Oudshoorn, 2011), were used to assess how opaque imputation models influence clinician trust and decision-making.
[…]
fig:IML_survey Refer to caption Refer to caption Refer to caption
@@ 2 Survey methodology
Data analysis.
We analyzed multiple-choice responses using descriptive statistics. For questions with responses on the Likert scale or categorical responses, we applied Factor Analysis for Mixed Data (FAMD) in R (Lê et al., 2008) to explore latent structures and identify response patterns. FAMD, the mixed data counterpart of Principal Component Analysis (PCA), accommodates both continuous and categorical variables, making it particularly well suited for questionnaire data (Josse and Husson, 2012). Since our study focuses on missing values, FAMD provides a natural framework where handling missingness is integrated into the analysis. To prevent bias from discarding incomplete responses, we employed the Expectation-Maximization (EM) algorithm to estimate principal components while maximizing the observed likelihood. This approach ensures that missing value estimation aligns with imputation, where minimizing the weighted least squares criterion or maximizing the observed likelihood produces equivalent results. To further analyze clinician groupings (Finding 3.1) based on AI/ML attitudes and incomplete data practices, we applied clustering to the FAMD results, allowing for a multivariate characterization of participants beyond isolated variables. Missing values in the survey answers, such as unanswered questions, were imputed using PCA for continuous variables and the most frequent category for categorical data. This process was implemented using the missMDA package (Josse and Husson, 2016), which relies on iterative Singular Value Decomposition (SVD) (Hastie et al., 2015), ensuring that imputed datasets preserve the same principal components as the original, maintaining the dataset’s underlying structure. Free-text responses were analyzed for central tendency and variability. Figure LABEL:fig:survey_example in the Appendix illustrates the full experimental setup.
Formal validation.
We applied systematic methods to formally validate our qualitative survey and enhance the reliability, credibility, and accuracy of the data collection (Rosellini and Brown, 2021). First, we targeted domain experts to ensure the survey’s relevance, clarity, and appropriateness. Before launching, we conducted a pilot study (Shakir and ur Rahman, 2022) with five independent trauma clinicians, incorporating their feedback to refine the clarity of the question and the response options. In Section 4, we applied triangulation by comparing our findings with other data sources, such as quantitative studies, and literature. Finally, to ensure consistency in qualitative analysis, multiple researchers independently reviewed the coded responses.
3 Best practices for IML with missing values at test time
This section presents findings aligned with the three main objectives (Figure LABEL:fig:objectives): clustering participants by their attitudes toward AI/ML in healthcare, summarizing clinician interactions with current IML models for handling test-time missingness, and outlining requirements for future IML research.
3.1 Finding 1: Clustering based on AI/ML attitudes and missing values practices
We received 55 survey responses, including 20 fully completed and 33 partially completed, all of which were analyzed, resulting in varying answer counts per question. The average participant age was 41.2 years, with a gender distribution of 22 females and 33 males. The majority of respondents were clinical doctors (83%), primarily in Anesthesiology/Intensive Care (84%), and 90% held at least a medical degree. Participant demographics are provided in Table LABEL:tab:stats_participants.